# YSO Jets Driven by Magnetic Pressure Generated through Stellar Magnetosphere–Disk Interaction

### v1

@@ 6 Conclusions
    During the early stages of the simulation, the stellar magnetosphere opens up, forming three distinct magnetic regions: (1) a polar region of open stellar field lines extending outward from the star; (2) a lower-latitude region of open stellar field lines threading into the disk, with field directions opposite to those in the polar region; and (3) an intermediate zone between the two, where oppositely directed poloidal magnetic fields coexist, a crucial region in the jet-launching process. In the inner disk, the open stellar field overwhelms the native disk field, facilitating the launch of a magnetically driven disk wind.
@@
    The jet is driven by toroidal magnetic pressure through a “load–fire–reload” cyclic process primarily involving magnetic field lines that connect the star and the vertically elevated disk surface. In the “load” stage, differential rotation along a magnetic field line between the star and the disk transfers rotational energy into the magnetic field, amplifying the toroidal component along field lines. This twisting is especially pronounced in a layer along the surfaces of the stellar magnetosphere and the disk, where it builds up a strong toroidal field. As the toroidal field strengthens in that layer, it generates a current in the cylindrical-$\hat{R}$ direction. This current, in turn, produces a vertical Lorentz force that propels the gas upward: the “firing stage” of the cycle. The release of toroidal magnetic energy drives the gas to expand in the $\hat{z}$ direction, launching the jet. Eventually, magnetic reconnection along the jet-launching field lines resets the field configuration, returning the magnetic field lines to the “two-legged” configuration, anchored at both the star and disk, ensuring that the system is continually “reloaded” for subsequent episodes of jet-launching.
@@
    Compared to the disk-only model presented in Y. Tu et al. (2025a), both the jet and disk wind are stronger in the stellar magnetosphere model, exhibiting higher velocities and greater mass fluxes. The disk wind is stronger because the partially opened stellar magnetosphere provides additional open magnetic flux to the disk, which increases the field strength in the inner disk and facilitates a stronger disk wind. The rotating star provides an additional power source to power the jet, which, together with the stronger open magnetic field in the polar cavity, helps drive a stronger jet in the magnetosphere model.
@@
    The jet tends to be bipolar in the magnetosphere model, in contrast to the disk-only setup, where the jet tends to be unipolar. This difference arises primarily from the geometry of the global magnetic field. In the disk-only case, the magnetic field structure typically facilitates the opening of a jet-launching cavity in only one hemisphere, while the other hemisphere is more likely to be choked by the disk wind. However, when a rotating stellar magnetosphere is included, the toroidal magnetic field reverses signs in both hemispheres, enabling the formation of polar cavities on both sides of the midplane. Consequently, jets driven by the magnetosphere-disk interaction are more likely to be bipolar and symmetric.
YT acknowledges support from the interdisciplinary fellowship at UVA. CYH acknowledges support from an ALMA Student Observing Support (SOS) from NRAO. We acknowledge computing resources from UVA research computing (RIVANNA), NASA High-Performance Computing, and NSF’s ACCESS computing allocation AST200032 and PHY250098. ZYL, YT, and CYH are supported in part by NASA 80NSSC20K0533, NSF AST-2307199, JWST-GO-02104.002-A, and the Virginia Institute of Theoretical Astronomy (VITA). Z.Z. acknowledges support from NSF award 2408207, 2429732 and NASA grant 80NSSC24K1285. {contribution} YT led the modeling, analysis, interpretation, and writing of the paper. ZYL and Z.Z. contributed to the analysis and interpretation, and oversaw the project’s progress. HX and CYH contributed to the interpretation.

### v2

@@ VI Conclusions
    Our 3D simulation reproduces the essential load–fire–reload magnetic cycle identified in earlier 2D studies (e.g., Zanni and Ferreira 2013) but extends it in several crucial ways. In the absence of ad hoc magnetic diffusivity or viscosity, the system develops a magnetically elevated, MRI-active inner disk that feeds mass onto open, star-anchored field lines via localized reconnection events. The third (azimuthal) dimension enables the cycle of loading, firing, and reloading to occur rapidly and asynchronously across azimuth—concurrently but without global coherence—producing a quasi-steady, continuously replenished jet rather than the large, periodic ejections typical of axisymmetric models. Thus, a fast, magnetically dominated jet can arise self-consistently from turbulent disk–magnetosphere coupling in ideal MHD.
@@
    The physical picture that emerges is that of an accretion-fed, star-anchored Poynting jet operating in the low-density polar cavity. Star-anchored field lines extract rotational energy from the star and inner disk, periodically reconnecting with the magnetized disk and its elevated atmosphere to reload mass and magnetic stress. Each field line cycles between a two-footpoint phase, where mass and twist accumulate, and a one-footpoint phase, where magnetic pressure accelerates plasma outward. Differential rotation regenerates the toroidal field, and asynchronous reconnection across azimuth maintains a continuous, large-scale jet. The jet in this scenario is lightly mass-loaded, as in the classical magneto-centrifugal model, but accelerated by magnetic pressure gradients, as in the magnetic-tower model. Its formation is facilitated by the low-density polar cavity kept open by star-anchored open field lines.
@@
    The global outflow consists of two coupled components: a fast, lightly loaded Poynting jet confined to the polar cavity and a slower, denser disk wind. The jet is not the densest axial part of the disk wind, which, in our simulation, is launched through the classical magnetic-tower mechanism. The denser disk wind dominates the system’s mass loss and provides a confining sheath for the jet, while the jet itself contains a significant toroidal magnetic field that contributes to self-collimation.
@@
    The stellar magnetosphere provides the star-anchored open field lines that maintain the bipolar cavities (one in each hemisphere), ensuring bipolar jets and preventing choking by turbulent disk material. In contrast to disk-only models that often produce one-sided or intermittent outflows, the inclusion of the rotating stellar field yields a persistent, bipolar structure in which the jet extracts energy and angular momentum from both the star and the disk. The appearance of similar fast, magnetically dominated jets across simulations with different microphysical assumptions—ranging from magnetically diffusive 2D (e.g., Ustyugova et al. 2006) to ideal MHD 3D—indicates that magnetic-pressure acceleration of the accretion-fed low-density material in the polar cavity to high speed through disk–magnetosphere interaction is a robust mechanism for jet production in young stellar systems. Higher-resolution simulations are desirable to firm up this conclusion.
We thank the referee for detailed and constructive comments that significantly improved the paper. YT acknowledges support from the interdisciplinary fellowship at UVA. CYH acknowledges support from an ALMA Student Observing Support (SOS) from NRAO. We acknowledge computing resources from UVA research computing (RIVANNA), NASA High-Performance Computing, and NSF’s ACCESS computing allocation AST200032 and PHY250098. ZYL, YT, and CYH are supported in part by NASA 80NSSC20K0533, NSF AST-2307199, JWST-GO-02104.002-A, and the Virginia Institute of Theoretical Astronomy (VITA). $Z.$ acknowledges support from NSF award 2408207, 2429732 and NASA grant 80NSSC24K1285.
{contribution}
YT led the modeling, analysis, interpretation, and writing of the paper. ZYL and Z.Z. contributed to the analysis and interpretation, and oversaw the project’s progress. HX and CYH contributed to the interpretation.